OpenStudy (anonymous):

Find the values of x in the following equation: |-5x + -2| = 12

OpenStudy (anonymous):

\[5x+2=12\] or \[5x+2=-12\] are your two equations.

OpenStudy (anonymous):

Find the values of x in the following equation: |-5x + -2| = 12

OpenStudy (anonymous):

He's explaining to you how to solve the problem. The absolute value is why you need to evaluate for both 12 and -12.

jhonyy9 (jhonyy9):

5x+2=12 and 5x+2=-12 5x=10 5x=-14 x=2 x= -14/5